# DKiS: Decay weight invertible image steganography with private key

### 30 Nov 2023 v1

@@ 5 Application
The integration of a private key significantly enhances the security of image hiding techniques, particularly for methods that are publicly accessible. For example, the use of a private key enables our method to address challenges that traditional image hiding approaches cannot solve. For instance, consider the application in photo verification. As illustrated in Fig. 9, camera manufacturers can embed a secret verification image into photos using a private key known only to them. If verification is required, users can submit the photo to the manufacturer. Successful extraction of the secret verification image by the manufacturer will confirm that the photo was taken with their product and has not been modified. This approach could become a universal standard among camera manufacturers, with each setting their unique private key. In the event of a key leak, the impact would be limited to the compromised manufacturer, while the integrity of other cameras remains unaffected.
@@ 6 Conclusion
In this paper, we introduced DKiS, a pioneering approach in the realm of image steganography, distinguished by its integration of a private key mechanism. Central to DKiS is the novel concept of decay weight, which strategically controls the transfer of information from the secret pipeline to the host pipeline. This is based on the insight that the relevance of information for the host pipeline diminishes progressively along the secret pipeline. The implementation of decay weight has notably enhanced the overall performance of the method.
A pivotal feature of DKiS is the incorporation of a private key into the steganographic process. This addition effectively addresses the longstanding challenge of balancing widespread accessibility with robust security in image steganography. As a result, DKiS emerges as a viable candidate for standardization in applications like image verification, where both the publicity and security of steganographic methods are paramount.
Our extensive experimental evaluations, conducted under conditions that include rounding errors, affirm the satisfactory performance of DKiS. The results demonstrate not only the method’s efficacy in securely hiding images but also its potential as a versatile tool in a wide range of practical scenarios. DKiS, therefore, stands as a significant contribution to the field of image steganography, bridging the gap between public accessibility and stringent security requirements.

### 18 Jan 2024 v2

@@ 5 Application
The integration of a private key significantly enhances the security of image hiding techniques, particularly for methods that are publicly accessible or likely attacked. For example, the use of the private key enables our method to address challenges that traditional image hiding approaches cannot solve. For instance, consider the application in photo verification. As illustrated in Fig. 11, camera manufacturers can embed a secret verification image into photos using a private key known only to them. If verification is required, users can submit the photo to the manufacturer. Successful extraction of the secret verification image by the manufacturer will confirm that the photo was taken with their product and has not been modified. This approach could become a universal standard among camera manufacturers, with each setting their unique private key. In the event of a key leaking, the impact would be limited to the compromised manufacturer, while the integrity of other cameras remains unaffected.
The application of DKiS in differentiating real photos/videos from AI-generated content is especially pertinent given the rapid advancements in AI imagery. Its capability to embed secret watermarks in sensitive and critical media, such as CCTV footage or press photographs, enhances security and ensures the authenticity of these sources. In an era where digital authenticity is increasingly challenged, DKiS offers a much-needed solution to uphold the integrity of digital media.
@@ 6 Conclusion
In this paper, we introduced DKiS, a pioneering approach in the realm of image steganography, distinguished by its integration of the private key mechanism. Central to DKiS is the novel concept of decay weight, which strategically controls the transfer of information from the secret pipeline to the host pipeline. This is based on the insight that the relevance of information for the host pipeline diminishes progressively along the secret pipeline. The implementation of decay weight has notably enhanced the overall performance of the method.
A pivotal feature of DKiS is the incorporation of a private key into the steganographic process. This addition effectively addresses the longstanding challenges of balancing widespread accessibility with robust security and vulnerable to attacks in image steganography. As a result, DKiS emerges as a viable candidate for standardization in applications like image verification, where both the publicity and security of steganographic methods are paramount.
Our extensive experimental evaluations, conducted under conditions that include rounding errors, affirm the satisfactory performance of DKiS. The attack simulation further proves its security. These results demonstrate not only the method’s efficacy in securely hiding images but also its potential as a versatile tool in a wide range of practical scenarios. DKiS, therefore, stands as a significant contribution to the field of image steganography, bridging the gap between public accessibility and stringent security requirements.